OpenStudy (anonymous):

What is the value of (2x)^-2/(3x)^-3?

OpenStudy (anonymous):

\[{ { {2x}^-2 } \over { {3x}^-3 } } = { 2 \over 3 } x^{-2 - 3 } = { 2 \over 3 } x^{-5}\]

OpenStudy (anonymous):

But my given choices are: A.9/4 B.2/3 C.-49

OpenStudy (anonymous):

27x/4

OpenStudy (lalaly):

\[\frac{4}{9}x^{-2--3}=\frac{4}{9}x\]